Solve by taking square roots:9p^2=4, and please show me all the work
Answer by stanbon(48568)

9p^2=4
p^2-(4/9)=0
(p-2/3)(p+2/3)=0
p=2/3 or p=-2/3
=========
Another way to do it:
9p^2=4
9p^2-4=0
(3p-2)(3p+2)=0
p=2/3 or p=-2/3
=
Cheers,
Stan H.